The tube A made from an aluminum alloy (E = 73 GPa has an outer diameter of 75 mm. It is used to support the rod B which is made from steel (E = 200 GPa) and has a 25 mm diameter. Calculate the minimum thickness required for the tube if the maximum deflection of the loaded end of the rod must be limited to .4 mm.
1
Expert's answer
2021-09-12T07:24:41-0400

Maximum load = "73× 10⁹ ×π(\dfrac{75×10^{-3}}4)²= 8.06×10⁷N"


Considering allowable stress = "200×10⁹\ Pa"


"\sigma = \dfrac PA ,\ A =\dfrac P{\sigma}"


"A =\dfrac{8.06×10^7}{200×10^9} = 4.03×10^{-4}\ m²"


Considering allowable extension = "4× 10^{-4}\ m"


"\delta =\dfrac{PL}{AE}, \ A = \dfrac{PL}{\delta E}"


"A= \dfrac{8.06× 10⁷ ×1}{4×10^{-4}×73×10^9}= 2.76\ m²"


Larger area governs,

"A =π(\dfrac{d²}4)"


"d =\sqrt{\dfrac{4A}π} =\sqrt{\dfrac{4× 2.76}{π}} =1.87\ m"